Consider the following relation.
y=−x−5
Find four points contained in the inverse. Express your values as an integer or simplified fraction.
{( , ), ( , ), ( , ), ( , )}

Answer by MathLover1(20849)


y=-x-5
Find four points contained in the inverse. Express your values as an integer or simplified fraction.
first find the inverse
y=-x-5......swap x and y
x=-y-5..........solve for y
y=-x-5
so, inverse y'=-x-5 is same as given function
An arbitrary set of 4 points could be for y=-1, 0, 1,+3+
-1'=-x-5->x=-4
0'=-x-5->x=-5
1'=-x-5->x=-6
3'=-x-5->x=-8
{(+-4 , -1 ), (-5+ , 0 ), (-6+ , 1 ), ( -8 ,3 )}